# Through the Lens of Human-Human Collaboration: A Configurable Research Platform for Exploring Human-Agent Collaboration

### v1

@@ 4.2.2. Findings
All five participants successfully completed the assigned tasks in both scenarios, which indicates a functional baseline of system usability. Most participants stated the platform is “easy to start” (P1), “nicely designed” (P4), and “quite straightforward” (P5). They particularly valued the visualizations in the result analysis module. However, the walkthrough revealed five recurring themes of usability challenges, which directly informed a set of targeted revisions to the researcher interface.
@@
Participants initially reported a high cognitive load on the experiment selection page. Particularly, they stated that the experiments are “not easy to understand at the first glance ” (P3) and the “the goal and the research focus are jammed together” (P1) in dense experiment introductory text, leading to an unclear mental model of the process, including the experiment rules, what roles exist, or how collaboration is facilitated. Besides, there is no indication of the procedures for experiment setup and their order, which causes the participants to be confused about what they should do and how they should proceed. Participants suggested using “mini-games” (P1) or “illustrations” (P4) to provide a more easy-to-comprehend onboarding process. In response, we replaced the experiment introductory text with a structured overview for each paradigm and included an illustration depicting the actors, interaction channels, and core task structure. This provides researchers with a better conceptual foundation before they engage with configuration settings. In addition, the procedures of the experiment setup workflow are naturally embedded with the horizontal scrolling feature.
Parameter Semantics and Previews for Interface Manipulation
Several participants reported struggling with interaction controls and experiment parameters, as the listing of experiment parameters looks “too overwhelming” (P1) during the session creation stage and the terminology was “computer science-y” (P1) and “vague” (P3). The struggle is attributed to the lack of a solid understanding of the experiments in the previous step, which makes it difficult to correctly interpret several experiment-specific parameters. Such difficulties were exemplified when the preliminary interface design listed too many controls without a structured organization, and they did not know what changes were made to the participant interface with those interaction controls. To address this, we implemented a live preview panel on the configuration screen. Adjusting an interaction control, such as communication channels or dashboard visibility, now immediately updates a preview of the corresponding participant interface module. In addition, we added more straightforward customization to the interaction controls; for instance, participants can select checkboxes to determine what information is and is not visible on the information dashboard.
Terminology and Information Architecture
The initial workflow enabled two customization concepts, “configuration templates” for experiment configurations and “session templates” for runtime configuration. Several participants confused these terms and reported “load template versus load session is very tricky” (P1). Terminology for interaction controls, such as “chat mode” and ”interaction variables,” also lacked clarity for participants.
In response, we simplified the model by removing the ”session templates.” Instead, the researchers now edit a single configuration template if they want to create custom experimental paradigms, and all the existing sessions are accessible by session ID. Further, we revised the terminology for all controls and parameters throughout the interface to match their functional scope without using specialized terminology. For instance, we use ”Interaction Control” to replace ”Interaction Variables,” use ”Private Chat” to replace ”Chat Mode,.”
Contextualized Configuration Across Tabs
Participants often needed to reference information from one stage while editing another. For instance, P2 shared she tends to “double-check” parameters at different tabs “before launching a session”. Participants may also need to adjust a parameter while checking the paradigm description or have new ideas for interaction controls after participant registration. However, the original tab-based navigation forced participants to switch between full-screen views, making it difficult to reference information from one stage while editing another. In response, we implemented a horizontal navigation functionality that allows users to seamlessly move between configuration tabs and experiment descriptions without losing context.
Results Inspection and Progressive Disclosure
Participants completed result analysis tasks with a smoother experience and higher satisfaction, as they described the plots and summary statistics as “easy to interpret” (P1) and “helpful for analyze” (P3). One participant’s complaint, which was also brought up in the session parameters page, was information density. Presenting all metrics as well as configurations at once makes it difficult for researchers to search for the relevant information they need. To improve this, we redesigned the results and configuration pages using a hierarchical structure with collapsible sections. This allows researchers to view a high-level summary by default and progressively disclose details as needed, improving information navigation and reducing information overload. This design was also implemented in the configuration pages.
@@
In summary, the cognitive walkthrough identified critical needs for a clearer conceptual entry point to each paradigm, a direct and visible connection between controls and their visible effects, simpler terminology, and smoother context maintenance. We revised the researcher interface design as well as the complete researcher workflow based on the feedback from the cognitive walkthrough study, which resulted in the final researcher interface presented in Section 3.2.
@@ 5. Discussion
We position this work as a methodological contribution to the HCI and CSCW communities. In an era of rapid technical advances in LLM agents, our platform provides a critical bridge between conventional human-human collaboration research, which is often facilitated through controlled human subject experiments, as well as the technical research frontier of LLM agents, which focuses on multi-agent frameworks and large-scale agent simulations.
5.1. An Open Platform for Controlled and Reproducible Human-Agent Interaction Experiments
Current research on LLM agents largely falls into two categories: multi-agent frameworks and social simulacra, both of which focus on demonstrating and advancing agents’ autonomous capabilities in complex environments. Examples of multi-agent frameworks include Agent-as-a-judge (Zheng et al., 2023a; Zhuge et al., 2024; Gu et al., 2024; Zhu et al., 2023), which asked agents to role-play human judges for decision-making, as well as GUI agents (Chen et al., 2025b; Zhou et al., 2023; Wang et al., 2025), where agents simulate personalized human interface-browsing behaviors given a dedicated persona profile. Social simulcara were proposed by HCI and social science researchers to form a virtual society and examine societal dynamics as LLM agents “live” in the virtual world (Park et al., 2023, 2024; Argyle et al., 2023; Aher et al., 2023; Almutairi et al., 2025; Chen et al., 2023). These efforts push the technical advances forward as agents demonstrate increasingly stronger autonomous task-solving, programming, and tool-using capacities in realistic settings.
Nevertheless, these systems often lack the rigorous controls necessary for human-in-the-loop experiments and do not focus on the nuances of human-agent interaction. The understanding of how the promises of these agents can be translated into practical benefits for humans that lead to a better joint outcome with agents lags behind technical innovations and has been overlooked by the technical community. Conversely, recent human–AI collaboration studies call for controlled, interaction-focused evaluation beyond aggregate task performance scores (Seeber et al., 2020; Zhang et al., 2023). Conventional HCI experiments primarily target computer-mediated human subject studies with careful task designs, variable controls, and analysis for collaborative research, such as DayTrader (Bos et al., 2002), Shape Factory (Bos et al., 2004), and NeoCITIES (McNeese et al., 2005; Schelble et al., 2022) Despite offering rigorous control over interaction variables, they are often costly to adapt for use and reproduce in the context of human-agent collaboration.
@@ 5.1. An Open Platform for Controlled and Reproducible Human-Agent Interaction Experiments
Our platform bridges the critical gap between these two research directions. It allows researchers to bring humans and LLM agents into the same experimental setup with mutual partnerships, while preserving the experimental paradigms, actions, states, communication channels, and equal partnership employed in classic HCI and CSCW research (Seeber et al., 2020; Zhang et al., 2023; Ju and Aral, 2025; Almutairi et al., 2025). In particular, a key mechanism enabling this contribution is the Agent Context Protocol that ensures consistent agent integration regardless of agent architectures. ACP ensures experimental parity by enforcing the perception and action spaces of agents to be similar to remote human collaborators, with respect to what the agent can perceive and do, how often it can act, and which parts of its state are visible. Subsequently, HCI researchers can isolate the effects of specific interaction controls and attribute changes to the specific variable, rather than to confounding factors in the agents’ design, behaviors, or the system. Our work lays a solid and reproducible foundation for building a cumulative understanding of human-agent collaboration.
5.2. Translating High-Level Guidelines to Testable Hypotheses for Human-Agent Collaboration
@@
A persistent challenge in HCI is how to translate high-level design guidelines into specific, testable hypotheses. Guidelines for human-AI interaction, for example, often recommend that systems should be explainable or that they should support efficient user control (Amershi et al., 2019; Liao et al., 2020). While valuable, it is often unclear how these principles should be implemented and what their behavioral impact is within a rich, interactive setting.
Our platform makes such guidelines empirically testable. Consider the guideline to ”make clear why the system did what it did” (Amershi et al., 2019). Our platform allows researchers to move beyond a simple ”explainability on/off” toggle and translate this as a set of experimental conditions. For example, one could compare an agent that provides no explanations, one that proactively offers a brief rationale for every trade offer, and one that provides explanations only upon request. Our platform logs the downstream effects on collaborative dynamics to examine how different conditions vary.
In addition, classic CSCW research has well-developed accounts of media effects (Straus, 1997), common ground (Clark and Brennan, 1991), awareness (Gutwin and Greenberg, 2002), and interdependence in distributed work (Olson and Olson, 2000), where the majority of prior tests only involve humans. Our platform can faithfully re-implement classic experimental paradigms while integrating LLM agents analogous to remote human participants. In particular, our method opens a shared research agenda, which allows HCI researchers to systematically investigate which collaboration principles remain valid when partners are LLM agents, which ones need revision, and what new patterns emerge. For example: when do richer channels (text + structured signals) actually improve grounding (Clark and Brennan, 1991) versus adding noise as media become more complex (Straus, 1997; Bos et al., 2002); when does higher agent initiative speed progress (Allen et al., 1999), and when does it crowd out human agency (Tapal et al., 2017; Parasuraman and Riley, 1997)?
Answering these questions requires controlled comparative analysis, reusable tasks, and detailed logs that connect interaction design to the collaborative behavior and outcomes between humans and LLM agents (Almaatouq et al., 2021; Chen et al., 2016). We view our method as a complement to existing study designs rather than replacements. For instance, researchers can use Wizard-of-Oz to probe early ideas and vocabulary (Kelley, 2018; Dahlbäck et al., 1993; Porcheron et al., 2021), use field work to study practice and organizational constraints (Rogers, 2011; Crabtree et al., 2013), and use the present platform when the goal is to make a claim about how a particular change in communication, awareness, initiative, or framing affects human collaboration with LLM agents.
5.3. Customizability and Generalizability for Interdisciplinary Research
The customizability and generalizability of our platform enable interdisciplinary researchers to benefit from human-agent collaboration experiments in enhancing agent capacities, developing agent-based systems, and designing human-agent interfaces (Almaatouq et al., 2021; Ju and Aral, 2025). For agent researchers, our platform can be viewed as a human-centered benchmark that provides far richer insights into the social and cognitive alignment of agent behaviors compared with agent-only frameworks. Instead of evaluating new agent architectures (e.g., with different memory or reasoning modules) only in agent-agent simulations (Park et al., 2023; Li et al., 2023; Packer et al., 2023; Yao et al., 2023; Wang et al., 2023b; Argyle et al., 2023; Aher et al., 2023), they can use our platform to test how these architectures perform when collaborating with humans on established tasks. For instance, findings from case studies in our paper can be useful not only in validating established human collaboration findings but also provide concrete design insights for agent researchers, particularly with respect to the differences (e.g., reasoning, planning, communicating) identified between humans and agents in equal collaboration scenarios (Seeber et al., 2020).
For social scientists, the platform offers a novel tool to test theories of group dynamics. One could, for example, program agents to consistently adopt specific negotiation strategies (e.g., tit-for-tat, always defect) and study how human teams react, providing a level of behavioral control that is impossible to achieve with human confederates (Axelrod, 1984).
For HCI researchers, in addition to key contributions presented in this paper, the platform also supports the creation of domain-specific ”scenario packs” that move beyond abstract tasks to model real-world collaborative workflows, while still maintaining experimental control. Despite classic experimental paradigms offering robust insights into media effects, awareness, and interdependence, the relatively simple task formation leaves open questions about how these findings generalize to real-world domains with different roles, norms, and risks (Olson and Olson, 2000; Gutwin and Greenberg, 2002). In response, our platform is designed for multi-human, multi-agent collaboration, with different roles, states, tasks, and constraints assigned to each participant. This design aligns with emerging domain studies that pair LLM agents with humans in situated workflows (e.g., LLM conversational agent to support patient-provider communication and daily health check-ins (Yang et al., 2024; Wu et al., 2025)).
To facilitate this broader aim, we are committed to the open-science practices and invite the community to co-develop reusable “scenario packs” (tasks, role files, UI variants), agent context templates, and anonymized log schemas. We provide open-source code and configuration profiles created by our teams. Our release follows the HCI toolkit guidance that evaluates contributions not only by controlled studies but also by adoption, extensibility, and evidence embodied in the software itself (Ledo et al., 2018). By standardizing the configuration of scenarios, prompts, agent-context schema, and visual elements, and encouraging artifact sharing, we aim to help the CHI and CSCW community move toward comparable, extensible studies of human–agent collaboration that everyone can reproduce, critique, and build upon.
@@ 6. Future Work and Limitations
Our work provides a platform for a new kind of controlled, reproducible research in human-agent collaboration. However, we acknowledge several limitations that also define a clear agenda for future work.
First, a primary limitation lies in the fidelity of the LLM agents. While our prompt engineering created plausible collaborators for the experimental task, there are plenty of opportunities to improve agents’ cognitive behaviors. Current LLM agents might be able to mimic behavioral patterns of humans, but they may not possess the deeper social and cognitive groundings that underpin human cognitive and human collaboration. This limitation constrains the direct generalizability of findings about trust and social dynamics from our platform to human-human collaboration research.
@@ 7. Conclusion
This paper addresses a critical gap between the rapid development of LLM agents and the need for controlled, reproducible methods to study human-agent collaboration. We presented an open, configurable research platform that bridges this gap by enabling researchers to re-implement classic CSCW experiments with LLM agents as analogous to remote human collaborators. Our system provides theory-driven interaction controls and a novel Agent Context Protocol to ensure experimental effectiveness and reproducibility. Our platform evaluation, conducted through two case studies involving the re-implementation of the Shape Factory experiment and a participatory cognitive walkthrough, demonstrates the platform’s effectiveness and usability in supporting controlled experiments and serving the research community. Our work opens up a broad avenue of human-agent collaboration research towards a systematic, evidence-based science.
@@ A.3. Shared Mental Model
A.4. Presence (Short et al., 1976)
@@ Appendix B Experiment Configuration Language
Appendix C Examples of Adapted Participant Interface

### v2

@@ 4.3.2. Findings
All five participants successfully completed both configuration and analysis tasks in both scenarios within the 45-minute session. This result provides an empirical baseline for the platform’s learning curve that new users can acquire the necessary mental model to operate the system’s GUI-based configuration layer in under an hour. Most participants stated the platform is “easy to start” (P1), “nicely designed” (P4), and “quite straightforward” (P5). They particularly valued the visualizations in the result analysis module. However, the walkthrough revealed four recurring themes of usability challenges, which directly informed a set of targeted revisions to the researcher interface.
@@
Participants initially reported high cognitive load on the experiment selection page. P3 stated that the experiments are “not easy to understand at the first glance ” while P1 noted that “the goal and the research focus are jammed together” in the dense introductory text. This difficulty resulted in an unclear mental model of the process; for instance, participants struggled to distinguish between ”experiment rules” (what the subjects do) and ”configuration options” (what the researcher does). Participants suggested using “mini-games” (P1) or “illustrations” (P4) to make onboarding easier to follow.
Parameter Semantics and Feedback
Several participants struggled with interaction controls and experiment parameters due to the lack of visible feedback during configuration. P1 described the parameters looks “too overwhelming” and P3 critiqued the terminology as “computer science-y” (P1) and “vague”. The struggle was attributed to a ”Gulf of Execution:” participants could not visualize how changing a parameter (e.g., ”Communication Level”) would affect the participant interface and interactions. Without immediate feedback, participants hesitated to modify defaults.
Information Architecture and Context Switching
The initial workflow introduced confusion between “configuration templates” (for experiment configurations) and “session templates” (for runtime configuration). P1 described that “load template versus load session is very tricky” (P1), which caused them to be confused about where to save the settings. Furthermore, P2 shared she tends to “double-check” parameters at different tabs “before launching a session”. The original tab-based navigation forced participants to switch between full-screen views, making it difficult to reference information from one stage (e.g., Role Assignment) while editing another (e.g., Interaction Controls).
Results Inspection
While participants praised the result analysis tools as “easy to interpret” (P1) and “helpful for analyze” (P3), they complaint about the information density. The interface presented all metrics as well as configurations at once, making it difficult for researchers to search for the relevant information they need.
4.3.3. Iterative Design Refinements
Based on these findings, we implemented a series of refinements to the final Researcher Interface (Section 3.2.2). These changes were designed to lower the learning curve by addressing the specific friction points identified above.
Scaffolding for Onboarding
To address the cognitive load in paradigm comprehension, we replaced the experiment introductory text with a structured overview for each paradigm and included an illustration depicting the actors, interaction channels, and core task structure. These visual summaries provide researchers a conceptual anchor for the setup process before configuring experiments.
Live Preview Panel
To bridge the gap between parameter configuration and interface manipulation, we implemented a real-time Live Preview. Adjusting an interaction control (e.g., communication channels or dashboard visibility), now immediately updates a preview of the corresponding participant interface module. We also added more straightforward customization to the interaction controls; for instance, participants can select check boxes to determine what information should or should not be visible on the information dashboard. Thus, the direct manipulation feedback loop clarifies parameter semantics without requiring technical terminology.
Streamlined Information Architecture
We simplified the terminology (e.g., renaming ”Chat Mode” to ”Private Chat”) and removed the complex ”Session Template” concept in favor of a unified configuration file. Furthermore, we replaced the tab views with a Horizontal Navigation bar and hierarchical menus. This allows researchers to maintain context while scrolling between configuration steps, while the interface progressively discloses more information as needed. This design was also implemented in the configuration pages.
@@ 4.3.3. Iterative Design Refinements
In summary, the cognitive walkthrough identified specific user needs for a clearer conceptual entry point, a direct feedback loop between backend controls and visible front-end effects, and smoother context maintenance during configuration. We integrated these insights to redesign the researcher interface and workflow.
@@ 5. Discussion
We position this work as a methodological contribution to the HCI and CSCW communities. By enabling rigorous variable control within a modern agent-integration environment, we provide a pathway for enabling human-agent interaction studies reproducible and analyzable in controlled settings, as well as moving from ad-hoc agent evaluation to systematic interactive analysis.
5.1. Bridging Agent Autonomy and Experimental Control
Researchers have identified novel capacities that enable LLM agents to think and behave like humans through role-playing divergent human persona characteristics. The rising research interest in multi-agent interaction platforms (Wu et al., 2024; Chen et al., 2023) and social simulacra (Park et al., 2023; Argyle et al., 2023) pushes the boundaries of what agents can do in complex environments. Specifically, these efforts show that agents can role-play as human judges (Zheng et al., 2023b; Zhuge et al., 2024; Chen et al., 2025c), simulate human web-browsing behaviors (Chen et al., 2025b; Zhou et al., 2023; Wang et al., 2025), and act as human participant surrogates in social experiments (Park et al., 2023, 2024; Argyle et al., 2023). However, as recent critiques note (Seeber et al., 2020; Zhang et al., 2023), these technical advances often outpace our understanding of how LLM agents should interact and collaborate with humans.
@@ 5.1. Bridging Agent Autonomy and Experimental Control
As LLM agents demonstrate promising potential to interact with humans just like remote collaborators, the systematic evaluation of human-agent interactions requires controls and grounding in established findings of remote human collaboration. Our research platform addresses this problem by optimizing the experimental parity between humans and LLM agents and prioritizing the experimental controls. Through the Agent Context Protocol (ACP), our platform enforces agents operate under the same perceptual and temporal constraints as remote human collaborators, while preserving the experimental paradigms, actions, states, communication channels, and equal partnership employed in classic HCI and CSCW research (Seeber et al., 2020; Zhang et al., 2023; Ju and Aral, 2025; Almutairi et al., 2025). Our platform allows HCI researchers to isolate the effects of specific interaction controls and attribute changes to the specific variable, rather than to confounding factors in the agents’ design, behaviors, or the system.
5.2. Translating Design Guidelines to Testable Hypotheses
@@
A persistent challenge in HCI is how to translate high-level design guidelines into empirically testable hypotheses. While guidelines for human-AI interaction often recommend that systems should be explainable or that they should support efficient user control (Amershi et al., 2019; Liao et al., 2020), verifying the impact on behaviors and perceptions of these principles requires controlled comparative analysis. Nevertheless, it is often unclear how these principles should be implemented and what their impact is within a rich, interactive setting. Our platform provides the technical scaffolding to convert these theoretical guidelines into configurable and empirically testable experimental conditions. For experimenting with ”explainability,” for instance, our platform allows researchers to compare an agent that provides no explanation, one that provides explanations upon request, and one that proactively offers a brief rationale for each action.
We view our method as a complement to existing study designs rather than a replacement. For instance, researchers can use Wizard-of-Oz to probe early ideas and vocabulary (Kelley, 2018; Dahlbäck et al., 1993; Porcheron et al., 2021), or use field work to study practice and organizational constraints (Rogers, 2011; Crabtree et al., 2013). Our platform opens a shared research agenda, which allows HCI researchers to systematically investigate which collaboration principles in human-human collaboration remain valid with LLM agent partners, which ones need revision, and what new patterns emerge.
5.3. Extensibility to Domain-Specific Contexts
Any system designed for extensibility faces an inherent trade-off between internal configurability (ease of use) and external customizability (flexibility). Frameworks for agent development like AutoGen (Wu et al., 2024) offer high flexibility for building complex agent behaviors but lack the experimental controls (e.g., logging, variable isolation) needed for immediate study deployment. Conversely, behavioral experiment platforms like Empirica (Almaatouq et al., 2021) offer robust experimental infrastructure but require significant engineering effort to integrate custom agent architectures, whereas game-based human-agent teaming platforms like Overcooked-AI (Carroll et al., 2019) and CREW (Zhang et al., 2024a) are designed for specific task scenarios. Our platform addresses this gap by combining a configurable researcher interface for interaction control manipulations with an open-sourced implementation that can be extended when projects require deeper customization.
While the experiment paradigms used in our evaluation are simplified, imaginary collaboration scenarios, the platform’s architecture offers a pathway toward broader application that incorporates domain-specific context. For instance, researchers can leverage the Experiment Configuration Language (ECL) to configure and adapt experimental mechanics (e.g., roles, resource distribution, information visibility) into the healthcare context to investigate how LLM agents, as ”AI nurses” in the out-patient team, can triage patients’ messages to the dedicated teams and coordinate the task hand-off. The structural flexibility distinguishes our platform from task-specific simulators like Overcooked (Carroll et al., 2019), allowing it to serve as a feasible research prototype for controlled studies of interaction variables with LLM agents. We believe that laying a solid foundation by isolating the effects of specific interaction variables in well-defined, controlled scenarios is a necessary precursor to studying complex, real-world domains. Future work will explore further lowering the barrier to entry for this expansion, potentially by integrating LLMs into the configuration workflow itself to translate natural language experiment descriptions into valid ECL templates.
@@ 6. Future Work and Limitations
Our work provides a platform for novel, controlled, reproducible research in human-agent collaboration. However, we acknowledge several limitations that also define a clear agenda for future work.
First, a primary limitation lies in the fidelity of the LLM agents. While our prompt engineering created plausible collaborators for the experimental task, there are plenty of opportunities to improve agents’ cognitive behaviors. Current LLM agents might be able to mimic behavioral patterns of humans, but they may lack the deeper social and cognitive grounding that underlies human collaboration. This limitation constrains the direct generalizability of our findings on trust and social dynamics to human–human settings.
@@ 7. Conclusion
This paper addresses a critical gap between the rapid development of LLM agents and the need for controlled, reproducible methods to study human-agent collaboration. We presented an open, configurable research platform that bridges this gap by enabling researchers to re-implement classic CSCW experiments with LLM agents analogous to remote human collaborators. Our system provides theory-driven interaction controls and a novel Agent Context Protocol to ensure experimental effectiveness and reproducibility. Our platform evaluation, conducted through two case studies involving the re-implementation of the Shape Factory experiment and a participatory cognitive walkthrough, demonstrates the platform’s effectiveness and usability in supporting controlled experiments and serving the research community. Our work opens up a broad avenue of human-agent collaboration research towards a systematic, evidence-based science.
Acknowledgements.
This work was supported in part by a Google Research Scholar Award, a Gift from Adobe, an NVIDIA Academic Hardware grant, an AnalytiXIN Faculty Fellowship, and the U.S. National Science Foundation under grants CNS-2426395 and CMMI-2326378. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the sponsors.
@@ A.1. Trust
Table 5. Questionnaire items measuring perceived trust. This table presents the trust-related survey statements used to evaluate participants’ perceptions of fairness, honesty, and negotiation behavior.
@@ A.2. Awareness
Table 6. Questionnaire items measuring workplace awareness. This table lists survey items designed to assess how well participants tracked others’ actions, presence, and activities during collaboration.
@@ A.3. Shared Mental Model
Table 7. Questionnaire items measuring shared mental models. The table contains items related to participants’ alignment of goals, understanding of tasks, and consistency of group knowledge.
A.4. Presence
@@
Table 8. Questionnaire items measuring social presence. The table provides items that capture the extent to which participants perceived the presence, attentiveness, and responsiveness of others in the distributed collaboration.
@@ Appendix B Experiment Configuration Language
Fig. 8. Example of experiment configuration for DayTrader. A sample script written in the experiment configuration language defines the structure of the DayTrader experiment, including roles, investment options, and payoff logic.
Appendix C Hidden Profile Experiment Results
Table 9 present the participants’ self-reported perception mesurements.
Dimension Mean SD t p Order Effect
Perceived Trust (Proactive) 4.31 $\uparrow$ 0.96 3.88 0.01** 1.31
Perceived Trust (Passive) 4.21 $\uparrow$ 1.37 2.50 0.04* 1.21
Workplace Awareness (Proactive) 5.04 $\uparrow$ 1.13 5.12 0.00** 2.04
Workplace Awareness (Passive) 5.31 $\uparrow$ 1.10 5.92 0.00** 2.31
Collaboration Effectiveness (Proactive) 5.08 $\uparrow$ 1.50 3.92 0.01** 2.08
Collaboration Effectiveness (Passive) 5.41 $\uparrow$ 1.13 6.03 0.00** 2.41
Social Presence (Proactive) 3.19 $\uparrow$ 1.22 0.43 0.68 0.19
Social Presence (Passive) 2.88 $\downarrow$ 1.19 -0.30 0.77 -0.12
Workload
 Mental Demand (Proactive) 12.62 $\uparrow$ 4.57 1.63 0.15 2.62
 Mental Demand (Passive) 13.12 $\uparrow$ 2.30 3.85 0.01** 3.12
 Physical Demand (Proactive) 3.62 $\downarrow$ 2.20 -8.20 0.00** -6.38
 Physical Demand (Passive) 5.00 $\downarrow$ 3.89 -3.63 0.01** -5.00
 Temporal Demand (Proactive) 5.12 $\downarrow$ 2.90 -4.75 0.00** -4.88
 Temporal Demand (Passive) 9.00 $\downarrow$ 5.21 -0.54 0.60 -1.00
 Performance (Proactive) 16.25 $\uparrow$ 3.85 4.60 0.00** 6.25
 Performance (Passive) 15.38 $\uparrow$ 4.93 3.09 0.02* 5.38
 Effort (Proactive) 12.12 $\uparrow$ 3.52 1.71 0.13 2.12
 Effort (Passive) 11.88 $\uparrow$ 3.52 1.51 0.18 1.88
 Frustration (Proactive) 5.62 $\downarrow$ 6.52 -1.90 0.10 -4.38
 Frustration (Passive) 6.88 $\downarrow$ 6.42 -1.38 0.21 -3.12
Anthropomorphism
 Fake–Natural (Proactive) 2.50 $\downarrow$ 1.41 -1.00 0.35 -0.50
 Fake–Natural (Passive) 2.50 $\downarrow$ 1.60 -0.88 0.41 -0.50
 Machine like–Human like (Proactive) 2.38 $\downarrow$ 1.41 -1.26 0.25 -0.62
 Machine like–Human like (Passive) 2.75 $\downarrow$ 1.91 -0.37 0.72 -0.25
 Unconscious–Conscious (Proactive) 3.00 $\downarrow$ 1.31 0.00 1.00 0.00
 Unconscious–Conscious (Passive) 2.62 $\downarrow$ 1.77 -0.60 0.57 -0.38
 Artificial–Lifelike (Proactive) 2.50 $\downarrow$ 1.31 -1.08 0.32 -0.50
 Artificial–Lifelike (Passive) 2.62 $\downarrow$ 1.77 -0.60 0.57 -0.38
Perceived Intelligence
 Incompetent–Competent (Proactive) 3.75 $\uparrow$ 1.39 1.53 0.17 0.75
 Incompetent–Competent (Passive) 3.25 $\uparrow$ 1.39 0.51 0.63 0.25
 Ignorant–Knowledgeable (Proactive) 3.62 $\uparrow$ 1.51 1.17 0.28 0.62
 Ignorant–Knowledgeable (Passive) 3.62 $\uparrow$ 1.30 1.36 0.22 0.62
 Irresponsible–Responsible (Proactive) 3.75 $\uparrow$ 1.39 1.53 0.17 0.75
 Irresponsible–Responsible (Passive) 3.88 $\uparrow$ 0.64 3.86 0.01** 0.88
 Unintelligent–Intelligent (Proactive) 3.88 $\uparrow$ 1.36 1.82 0.11 0.88
 Unintelligent–Intelligent (Passive) 3.62 $\uparrow$ 1.30 1.36 0.22 0.62
 Foolish–Sensible (Proactive) 3.88 $\uparrow$ 1.36 1.82 0.11 0.88
 Foolish–Sensible (Passive) 4.00 $\uparrow$ 0.76 3.74 0.01** 1.00
Table 9. Paired t-test results for participants’ perceptions in the Hidden Profile experiments. Green arrows ($\uparrow$) indicate mean values above the midpoint, while red arrows ($\downarrow$) indicate those below. Statistically significant $p$-values are bolded, where * denotes $(p\leq 0.05)$, and ** denotes $(p\leq 0.01)$.
Table 9. Paired t-test results for participants’ perceptions of distributed collaboration in Hidden Profile experiments. Columns list the measured dimension, mean, standard deviation, t value, p value, and order effect. Dimensions include perceived trust, workplace awareness, collaboration effectiveness, social presence, workload components (mental, physical, temporal demand, performance, effort, frustration), anthropomorphism, and perceived intelligence. Green arrows indicate mean values above the scale midpoint, and red arrows indicate values below it. Statistically significant p values are bolded (p ≤ .05 marked with *, p ≤ .01 with **).
Appendix D Examples of Adapted Participant Interface